y" -5y' +4y=0 ;y(0)=1 and y' (0)=1 .


1
Expert's answer
2021-02-04T14:32:01-0500

"y'' -5y' +4y=0 ; \space \space\space \space y(0)=1 \space\space \space and \space \space\space y' (0)=1 ."

Setup the auxilary equation

"\lambda^2 - 5\lambda +4 = 0 \to \lambda_1= 1 \space \space\space \space \lambda_2=4"

The general solution is found by

Usually problem like these will have the answer in the form "C_1e^a+C_2e^b..." where a and b  are the roots of the characteristic equation "e^{rt}"

y = "C_1e^{t} + C_2e^{4t}" "\to C_1+C_2 = 1"

"y' = C_1e^t+4C_2e^{4t} \to C_1+4C_2 = 1"

"C_1 = 1 \space\space\space\space C_2= 0"

The general solution is: "y = e^t"